What is the derivative of ln"( (x^3 - 1) \/ x^2 )"


1
Expert's answer
2021-06-11T10:27:35-0400
"y=\ln(\dfrac{x^3-1}{x^2})"

"x^3-1>0=>x>1"

Then


"y=\ln(\dfrac{x^3-1}{x^2})=\ln(x^3-1)-2\ln x"


"y'=\ln(\dfrac{x^3-1}{x^2})'=(\ln(x^3-1)-2\ln x)'"

"=\dfrac{3x^2}{x^3-1}-\dfrac{2}{x}=\dfrac{x^3+2}{x(x^3-1)}"